Consider the case of a one-dimensional harmonic oscillator on the x-axis. Determine the Hamiltonian of the system using the Lagrangian transformation, then determine the equation of motion!


1
Expert's answer
2021-06-30T11:47:09-0400

simple harmonic oscillator


Hamiltonian simple harmonic oscillator -


Hamiltonian of a system is defined as -


"H=\Sigma{p_i}{q_i}""-L"


in the above equation "{p_i}=Generalised \ momentum"

"q_{i}=" "Generalised\ velocity"

"L=Lagrangian"


now we know that "L=T-V......1)" , and motion is along x-axis .


"T=\dfrac{MX'^{2}}{2}"


"V=\dfrac{KX^{2}}{2}"


Now putting "T\" and "V" in equation 1 , we get =


"L=" "\dfrac{MX'^{2}}{2}" "-\dfrac{KX^{2}}{2}"



Now , similarly Hamiltonian equation will become -


"H=" "\Sigma{p_i}{q_i}-L"


"H=\Sigma{p_i}{q_i}" "-\dfrac{Mx'^{2}}{2}+\dfrac{Kx^{2}}{2}"


"P_{i}=" actually momentum of spring along X-axis .


"Q_{i}=" actually generalised velocity along X-axis.


From lagrangian "\dfrac {\partial l}{\partial x}=p_{x}" , "\dfrac {\partial l}{\partial q_{c}'}=P_{i}"


"\dfrac {\partial l}{\partial x'}=" "\dfrac{\partial( Mx'^{2}+Kx^{2})}{2}"



"\dfrac {\partial l}{\partial x'}=mx'"


"\implies" "p_{x}=mx'\implies" "x'=\dfrac{Px}{m}"


now substituting the value of "x'" in Hamiltonian equation , we get -


"H=P_x\dfrac{P_x}{m}-\dfrac{{(mx')}^{2}}{2m}+\dfrac{Kx^{2}}{2}"



"H=\dfrac{P_x^{2}}{m}-\dfrac{P_x^{2}}{2m}+\dfrac{Kx^{2}}{2}"



"H=" "\dfrac{P_x^{2}}{2m}" "+\dfrac{Kx^{2}}{2}"


The above equation is Hamiltonian equation for simple harmonic oscillator .